TFY4340 Mesoscopic Physics Spring 2010

Course information

Updated: January 29, 2010.

Course webpage:

http://web.phys.ntnu.no/~stovneng/TFY4340_2010/tfy4340.htm

Lecturer: Jon Andreas Støvneng Department of physics, NTNU, 7491 Trondheim Office: E5-130 Phone: 73 59 36 63 Email: jon.stovneng@ntnu.no

Timetable:

Day  Hours  Room  Activity
Monday 10.15 – 13.00 E5-103 or B21 Lectures
Tuesday 08.15 – 10.00 E5-103 or B21 Exercise

Comment: Room E5-103 is in the Theory section of Department of physics, on level 5, E corridor.

Syllabus:

• To a first approximation, lectured material and given exercises.

• A more precise definition will come later.

Books etc: Parts of the books by Heinzel and Datta will cover most of the syllabus.

Evaluation: The exam will be on May 12.

Department of physics, January 29, 2010. Jon Andreas Støvneng